OpenStudy (anonymous):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: |x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day? A. x = −9, x = 21 B. x = −6, x = 18 C. x = 6, x = −18 D. No solution

OpenStudy (jango_in_dtown):

|x-6|+14=38 implies |x-6|=38-14=24 therefore x-6=-24,24 implies x=-18,30 so min =-18 nd max=30

OpenStudy (anonymous):

but that's not one of the answers

OpenStudy (jango_in_dtown):

since its called cold day, no 30 will never be reached... right? So you may go for option D

OpenStudy (ybarrap):

Since $$ |x-6|=38-14=24 $$ then $$ \sqrt{(x-6)^2}=24\\ (x-6)^2=24^2\\ x^2-12x+36=24^2\\ x^2-12x+36-24^2=0\\ \text{Using quadratic formula:}\\ x=-18,30 $$ This gives you the range of values that x can be